[Brahms-l] Shift report 20040209 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Mon Feb 09 2004 - 23:39:29 EST
------------------------------------------------------------
Date: Mon Feb  9 23:39:28 2004, Shift: 16:00-24:00
Supervisor: Ionut ARSENE, with: 
Log book#15 pages 51 - 51

Shift Summary:
------------------------------------------------------------
1. Useful beam for about 4 hours.

2. Data runs: 10234, 10235, 10236
 I finished the A2442 setting with FS at 2.3 degrees.
 In MRS we still need something like 2M trig3 events.
   The 10225 run is not good because we had problems with the DAQ (see E-log).
  10238 and 10239 are pedestal runs.

3. No detector problems observed. Everything worked smooth.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Feb 9 23:39:48 2004

This archive was generated by hypermail 2.1.8 : Mon Feb 09 2004 - 23:40:10 EST